A 500-ml aqueous solution initially contained 0.001M HZX. During equilibrium, ZX- ions were measured and the concentration was 5 x 10-13 M. Determine the chemical equilibrium constant expression and the chemical equilibrium constant.



1
Expert's answer
2021-03-29T06:04:03-0400

Chemical equilibrium constant expression

H2O + HZX -> H3ZXO

= H3ZXO/ (H2O)(HZX)

Chemical equilibrium constant

= K = H3ZXO/ (H2O)(HZX)

= 1 + 0.001 = (1.001/1×1) × 5×10^-13

K = 5×10^-13